40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 15 01:41:02 2013, Manasa, Yuta, Summary, Alignment, IFO aligned and ready for PRMI locking 
    Reply  Fri Feb 15 01:51:43 2013, Jenne, Summary, Alignment, IFO aligned and ready for PRMI locking 
Message ID: 8085     Entry time: Fri Feb 15 01:41:02 2013     Reply to this: 8086
Author: Manasa, Yuta 
Type: Summary 
Category: Alignment 
Subject: IFO aligned and ready for PRMI locking 

[Yuta, Manasa, Jenne, Jamie, Steve]

IFO aligned and ready for PRMI locking

Alignment procedure

0. Measured MC centering (off by 5mrad) before getting the doors off.

1. Got the TTs to 0.0 in pitch and yaw.

2. Using the MMTs, the beam was centered on the TTs.

3. TT1 was adjusted such that the incident beam was centered at PRM (with target).

4. TT2 was adjusted such that the beam passed through the center of BS (with target).

5. Centered the beam on PR2 by sliding it on the table.

6. Moved PR2 and tweaked TT2 to center the beam on ITMY and BS respectively.

7. Using TTs, we got the beam centered on ETMY while still checking the centering on ITMY.

8. ITMY was adjusted such that it retro-reflected at the BS.

9. ETMY was aligned to get a few bounces in the arm cavity.

10. Centered on ITMX by adjusting BS and then tweaked ITMX such that we retro-reflected at BS.

11. At this point we were able to see the MI fringes at the AS port.

12. Tweaked ITMX to obtain reflected MI fringes in front of MMT2.

13. By fine adjustments of the ITMs, we were able to get the reflected MI to go through the faraday  while still checking that we were retro-reflecting at the BS.

14. Tweaked the PRM, such that the PRM reflected beam which was already visible on the 'front face back face of faraday' camera went through the faraday and made fine adjustments to see it fringing with the reflected MI that was already aligned.

15. At this point we saw the REFL (flashing PRMI) coming out of vacuum unclipped and on the camera.

16. Started with alignment to get the AS beam out of vacuum. We tweaked OM1 and OM2 (steering mirrors in the ITMY chamber) to center the beams on OM4 and OM3 (steering mirrors in the BSC) respectively.

17. We then adjusted steering mirrors OM5 and OM6 (in the OMC chamber) such that the beam went unclipped out of vacuum.

18. Note that we took out the last steering mirror (on the AS table) in front of the AS camera, so that we can find the AS beam easily. This can be fixed after we pump down.

 

 

Tomorrow

  0. REFL still looks like an egg, but leave it .

  1. Align PRMI (no more in-vac!) .

  2. Align POP/REFL/AS cameras and PDs.

  3. Setup PRM/BS/ITMX/ITMY oplevs.

  4. Balance the coils on these mirrors.

  5. Lock PRMI.

ELOG V3.1.3-